OpenStudy (anonymous):

log12 (4) + 1/ 2log36 (2) + log36 (3) evaluate the expression

OpenStudy (anonymous):

\[\log_{12}(4)+\frac{ 1 }{ 2}\log_{36}(2) +\log_{36}(3)\] Is this your expression?

OpenStudy (anonymous):

no, 1 over (2log36 (2)+log 36 (3))

OpenStudy (amoodarya):

\[\log_{b^n}a^m=\frac{ m }{ n } \log_{b} a\]

OpenStudy (imer):

@kY_Tan :What you think the first step should be?

OpenStudy (anonymous):

log12 (4)+ 1/(log36 (4)+log36 (3)) log12 (4)+ 1/(log36 (12))

OpenStudy (tkhunny):

Where did we get a 4? \(\log(2) + log(3) = log(6)\)

OpenStudy (imer):

@tkhunny: you missed the "2" before \[\log_{36} (2)\]

OpenStudy (tkhunny):

Ah, that makes more sense.

OpenStudy (imer):

In proper terms;\[\log_{12}(4)+ \frac{1 }{ 2\log_{36} (2)+\log_{36} (3) }\]

OpenStudy (imer):

So what should be the next step @kY_Tan ?

OpenStudy (anonymous):

and then log12(4) +log36 (36)/log36 (12)?

OpenStudy (anonymous):

i tried for it many times but can't get the correct answer

OpenStudy (imer):

"log36 (36)/log36 (12)" is same 1/log36 (12), so you are not going anywhere.

OpenStudy (anonymous):

can't log36 (36)/log36 (12) be divided?

OpenStudy (imer):

We have "2" different bases, you need to make one of them have equal base to the other.

OpenStudy (imer):

It is divided but you still have "2" different bases, one with "12" and one with "36"

OpenStudy (anonymous):

so what i already done until now is correct?

OpenStudy (imer):

change the base to "12" by \[\frac{ \log_{12}(12) }{ \log_{12}(36) }\]

OpenStudy (anonymous):

so after changing the base to 12 1/(log12 (12)/log12 (36))?

OpenStudy (imer):

\[\log_{12} (4)+\frac{ 1 }{ \frac{ \log_{12}(12) }{ \log_{12}(36) } }\]

OpenStudy (imer):

But we know \[\log_{12} 12=1\]What would be the next step?

OpenStudy (imer):

@ kY_Tan : yes, you are correct!

OpenStudy (anonymous):

can log12 (4) + 1/(1/3)?? or just log12 (4) + log12 (36)?

OpenStudy (imer):

Don't simplify individual log function which do not provide discrete result.

OpenStudy (imer):

\[\log_{12} (4)+\log_{12} 36=\log_{12} (36*4)=\log_{12} (144)\]

OpenStudy (imer):

and sure we know; \[12^x=144\], what would be "x"?

OpenStudy (anonymous):

oh ok i got it 12^x=144 x=2

OpenStudy (imer):

Perfect! :)

OpenStudy (anonymous):

thanks for your guide

OpenStudy (imer):

My pleasure :)